# A Machine Learning and Finite Element Framework for Inverse Elliptic PDEs via Dirichlet-to-Neumann Mapping

### v1

@@ Abstract
Inverse problems for partial differential equations (PDEs) are crucial in numerous applications such as geophysics, biomedical imaging, and material science, where unknown physical properties must be inferred from indirect measurements. In this work, we address the inverse problem for elliptic PDEs by leveraging the Dirichlet-to-Neumann (DtN) map, which captures the relationship between boundary inputs and flux responses. Thus, this approach enables to solve the inverse problem that seeks the material properties inside the domain by utilizing the boundary data. Our framework employs an unsupervised machine learning algorithm that integrates a finite element method (FEM) in the inner loop for the forward problem, ensuring high accuracy. Moreover our approach is flexible to utilize partial observations of the boundary data, which is often the case in real-world scenarios. By incorporating carefully designed loss functions that accommodate discontinuities, the method refines coefficient reconstructions iteratively. This combined FEM and machine learning approach offers a robust, accurate solution strategy for a broad range of inverse problems, enabling improved estimation of critical parameters in applications from medical diagnostics to subsurface exploration.
@@
inverse problem, elliptic PDEs, EIT, variable coefficient, boundary condition mapping
@@ 1 Introduction
Inverse problems for partial differential equations (PDEs) are a fundamental area of study in applied and computational mathematics, with wide-ranging applications in fields such as geophysics [1], biomedical imaging [2, 3], and material science. These problems center on deducing unknown coefficients—often representing physical properties such as conductivity, diffusivity, or elasticity—from observed data [4, 5]. For instance, in medical imaging techniques such as Electrical Impedance Tomography (EIT), recovering these hidden material properties from indirect measurements enables visualization of conductivity distributions inside the human body, thereby enhancing diagnostic capabilities [6, 7, 8]. Similarly, in geophysical contexts, inferring subsurface properties such as permeability from surface measurements is critical for resource exploration and environmental monitoring [9, 10].
To address these challenges, we focus on the inverse problem for an elliptic equation with variable, potentially discontinuous coefficients. The main novelty of our approach is the utilization of the Dirichlet-to-Neumann (DtN) map—a mathematical construct that encapsulates the relationship between prescribed boundary inputs and their resulting flux responses [11]. This mapping framework is especially advantageous because it allows for the reconstruction of the interior characteristics of a domain using solely boundary measurements. In scenarios such as EIT or subsurface exploration, where sensors are predominantly deployed on the boundary, this capability eliminates the need for invasive internal data collection while still providing detailed insight into the material properties within the domain. See Figure 1 for the illustration.
@@
The proposed computational framework further distinguishes itself by integrating an unsupervised machine learning algorithm with finite element methods (FEM) used in the inner loop for solving the forward problem. This combination not only achieves high accuracy in modeling the physical processes but also enables iterative refinement of the coefficient reconstructions. The carefully designed loss functions, tailored to handle discontinuities effectively, enhance the overall fidelity of the inverse solution, making our approach robust and versatile. Ultimately, this framework offers a precise and reliable estimation of critical parameters, paving the way for significant advancements in applications ranging from non-invasive medical diagnostics to environmental and subsurface monitoring.
The rest of the paper is organized as follows. First, the governing system and the DtN mapping are introduced in Section 2. The overall computational framework including the loss functions is discussed in Section 3. Finally, in Section 4, several numerical experiments to illustrate the capabilities and the performance of our algorithm are presented.
@@ 2 Governing System
In this paper, we consider the following equation
@@
where $\Omega$ is a bounded domain in $\mathbb{R}^{2}$, $p:\overline{\Omega}\rightarrow\mathbb{R}$ is the solution function, $f$ is the source function, $g\in H^{1/2}(\partial\Omega)$ is the given Dirichlet boundary condition, and $k$ is a strictly positive coefficient defined in $\Omega$.
We consider an inverse problem associated with (1), where the goal is to reconstruct the coefficient $k$ from the boundary measurements. To achieve this, we employ deep learning techniques. A straightforward approach would be to use supervised learning, which requires access to exact coefficient values of $k({\mathbf{x}})$ for all ${\mathbf{x}}\in\Omega$. However, in most cases, obtaining complete knowledge of $k$ in $\Omega$ is infeasible. Thus, we propose an unsupervised learning framework that does not rely on exact coefficient values of $k({\mathbf{x}})$ but instead leverages available data to infer the coefficient.
@@ 2.1 Dirichlet-to-Neumann Map
In this subsection, we introduce the (Dirichlet-to-Neumann) DtN map, which is central to the reconstruction of $k$. This map is defined as follows:
@@
The reconstruction of the coefficient $k$ using the DtN map has been extensively studied in the context of Electrical Impedance Tomography (EIT) and the Calderón problem [11]. In particular, the case when $f=0$ has been comprehensively surveyed in [12]. Furthermore, additional studies have explored the uniqueness of $k$ under partially observed boundary data. Let $\Gamma\subset\partial\Omega$ be a non-empty open subset of the boundary. The following theorem proves the uniqueness of $k$.
Theorem A (Corollary 1.19 in [12])
@@
$\Lambda_{k_{1}}(g)=\Lambda_{k_{2}}(g)\text{ on }\Gamma\text{ for all }g\in H^{1/2}(\partial\Omega),\text{ supp }g\subset\Gamma.$
@@
Although the case $f\neq 0$ may seem easier because $f$ contains internal information, it remains an open problem. While the uniqueness of $k$ in the entire domain $\Omega$ is not guaranteed, its uniqueness on the boundary $\partial\Omega$ has been established [13].
Theorem B (Theorem 5.1.1 in [13])
@@
[…]
If $k_{1},k_{2}\in C(\Omega)$ are two coefficients that are analytic in $\Omega$ and $\Lambda_{k_{1}}=\Lambda_{k_{2}}$, then $k_{1}=k_{2}$ on $\partial\Omega$ and by Strong Maximum Principle [14, Chapter 2], we have $k_{1}=k_{2}$ in $\Omega$.
Let $\omega$ be either $\Gamma$ or $\partial\Omega$. We refer to the case $\omega=\Gamma$ as the partially observed case, and $\omega=\partial\Omega$ as the fully observed setting. Instead of directly using the DtN map, which is usually difficult to obtain in real-world scenarios, we utilize Cauchy data—pairs of Dirichlet and Neumann boundary measurements.
$\left\{(g,h_{g}):g\in H^{1/2}(\partial\Omega),\text{ supp }g\subset\omega,h_{g}=k\partial_{\nu}p_{g}|_{\omega}\right\}.$
These pairs serve as practical substitutes for the DtN map in our setup. Again, in practice, it is not feasible to train on all functions $g\in H^{1/2}(\partial\Omega)$ with supp $g\subset\omega$. We consider the following problem.
Problem
Given a sufficiently large finite collection of Cauchy data:
$\left\{(g_{i},h_{g_{i}})_{i=1,\dots,N}:g_{i}\in H^{1/2}(\partial\Omega),\text{supp }g_{i}\subset\omega,h_{g_{i}}=k\partial_{\nu}p_{g_{i}}|_{\omega}\right\},$ (2)
@@
Rather than directly using the values of $k$, we adopt an unsupervised learning framework where the model learns to reconstruct $k$ from Cauchy data data alone [15, 16]. Using this approach, we reconstruct $k$ based only on Cauchy data, which is particularly useful in practical setups where $k$ is not directly measurable.
@@ 3 Computational Algorithm
In this section, we present the numerical approach for solving Problem. First, in Subsection 3.1, we provide an overview of the proposed framework. We introduce the model for reconstructing the coefficient $k$ in Subsection 3.2. Next, in Subsection 3.3, we describe how to solve the forward problem using the model described in 3.2. Finally, in Subsection 3.4, we define the loss function used to train the model and describe the method for calculating its gradient to update the model parameters.
@@ 3.1 Overall Framework
Input $\bullet$ Domain points
$\overline{\Omega}$
$\bullet$ Neumann data
$h_{g}=k\partial_{\nu}p_{g}|_{\omega}$
Neural Network $\mathcal{N}$
$\bullet$ Multilayer Perceptron (MLP)
: Computation results: Data flow Loss Function
$\mathcal{L}\bigl{(}h_{g},\tilde{h}_{g}\bigr{)}=\bigl{\|}h_{g}(\cdot)-\tilde{h}_{g}(\cdot)\bigr{\|}^{2}_{L^{2}(\omega)}$
Output $\bullet$ Reconstruct Coefficients
$\{\tilde{k}({\mathbf{x}})\,|\,{\mathbf{x}}\!\!\in\!\!\overline{\Omega}\}$
Predicted Neumann
$\tilde{h}_{g}=\tilde{k}\,\partial_{\nu}\tilde{p}_{g}|_{\omega}$
Solving for the PDE (FEniCS) Input $\bullet$ Source function $f$
$\bullet$ Dirichlet condition $g$
$\tilde{p}_{g}({\mathbf{x}})=\sum_{j=1}^{M}p_{j}\,\phi_{j}({\mathbf{x}})$ Parameter UpdatesDeep LearningFinite Element Method
@@
Figure 2 illustrates the process of reconstructing the coefficient $k$ using a neural network and FEM. The framework proceeds as follows: First, the domain points $\overline{\Omega}$ are input to the neural network $\mathcal{N}$, which outputs the reconstruction $\tilde{k}$ of $k$. The neural network $\mathcal{N}$ is designed as a multilayer perceptron (MLP) that reconstructs $k$ from the input domain points. Next, the reconstructed coefficient $\tilde{k}$, along with the Dirichlet condition $g$ and source function $f$, is used to solve the (1) via FEM with the FEniCS package. The resulting predicted solution is denoted as $\tilde{p}_{g}$, which is represented as a linear combination of basis functions $\{\phi_{i}\}_{i=1}^{M}$ in the finite-dimensional subspace $V_{h}$. These basis functions and the finite-dimensional subspace are further detailed in Subsection 3.3. Using $\tilde{k}$ and $\tilde{p}_{g}$, the predicted Neumann data $\tilde{k}\partial_{\nu}\tilde{p}_{g}|_{\omega}$ is computed. This predicted Neumann data is then compared with the given data $h_{g}$ to calculate the loss function. Finally, the model parameters of the neural network $\mathcal{N}$ are updated to minimize the loss function. By iterating through this process, the neural network $\mathcal{N}$ learns to reconstruct $k$ more accurately.
@@ 3.2 Model for Reconstructing the Coefficient $k$
In this subsection, we propose a model for reconstructing the coefficient $k$. The model is based on a MLP architecture, which leverages the Universal Approximation Theorem [17] as its theoretical foundation. This theorem ensures that neural networks, including MLPs, can approximate any continuous function to a desired level of accuracy.
The neural network $\mathcal{N}$ takes the spatial variable ${\mathbf{x}}$ as input and reconstructs the coefficient $k$ as $\tilde{k}$. Specifically, the reconstruction is expressed as $\tilde{k}({\mathbf{x}})=\mathcal{N}({\mathbf{x}};\boldsymbol{\theta})$, where $\boldsymbol{\theta}$ represents the learnable parameters of the model, such as weights and biases. The network consists of three hidden layers, each containing 50 hidden units. To enhance the ability of the model to capture periodic patterns and their derivatives, we employ the sine function as the activation function in the hidden layers. This choice enables the model to efficiently represent complex natural signals with high accuracy.
3.3 Forward Problem
In this subsection, we briefly describe the forward problem for the solution $p$ of (1) by utilizing FEM. Here, we employ the linear continuous Galerkin method, and it is implemented by FEniCS [18]. The problem assumes a given Dirichlet condition $g$ and coefficient $k$, and the problem (1) is reformulated into a weak form by integration and then approximating the solution in a finite-dimensional space.
In particular, let $V_{h}$ be a finite-dimensional subspace of $H^{1}(\Omega)$, and assume $p_{h},v_{h}\in V_{h}$, with $p_{h}$ representing the approximate solution and $v_{h}$ being the test function used in the weak formulation. Using the basis functions $\{\phi_{i}\}_{i=1}^{M}$ of $V_{h}$, the weak form can be approximated as:
$\int_{\Omega}k\nabla p_{h}\cdot\nabla v_{h}\,{\rm d}{\mathbf{x}}=\int_{\Omega}fv_{h}\,{\rm d}{\mathbf{x}}.$
Assuming $p_{h}=\sum_{i=1}^{M}P_{i}\phi_{i}$ and substituting $v_{h}=\phi_{j}$, the resulting system for each $j$-th term is:
$\sum_{i=1}^{M}P_{i}\int_{\Omega}k\nabla\phi_{i}\cdot\nabla\phi_{j}\,{\rm d}{\mathbf{x}}=\int_{\Omega}f\phi_{j}\,{\rm d}{\mathbf{x}}.$

### v3

@@ Abstract
Inverse problems for Partial Differential Equations (PDEs) are crucial in numerous applications such as geophysics, biomedical imaging, and material science, where unknown physical properties must be inferred from indirect measurements. In this work, we present a new approach to solving the inverse problem for elliptic PDEs, using only boundary data. Our method leverages the Dirichlet-to-Neumann (DtN) map, which captures the relationship between boundary inputs and flux responses. This enables the reconstruction of the unknown physical properties within the domain from boundary measurements alone. Our framework employs a self-supervised machine learning algorithm that integrates a Finite Element Method (FEM) in the inner loop for the forward problem, ensuring high accuracy. Moreover, our approach illustrates its effectiveness in challenging scenarios with only partial boundary observations, which is often the case in real-world scenarios. In addition, the proposed algorithm effectively handles discontinuities by incorporating carefully designed loss functions. This combined FEM and machine learning approach offers a robust, accurate solution strategy for a broad range of inverse problems, enabling improved estimation of critical parameters in applications from medical diagnostics to subsurface exploration.
@@
FEM, elliptic PDEs, inverse problem, Dirichlet-to-Neumann map, partial boundary observation
journal: Journal of Computational and Applied Mathematics
@@ 1 Introduction
Inverse problems involve inferring unknown properties from indirect observations, in contrast to forward problems, where the governing equations and parameters are given and the solution is computed directly. Such problems often arise in science and engineering, particularly when reconstructing coefficients, sources, or initial conditions in governing equations. In particular, inverse problems for Partial Differential Equations (PDEs) are a fundamental area of study in applied and computational mathematics [Isakov06, Kirsch11], with wide-ranging applications in fields such as geophysics [Duchateau97], biomedical imaging [Kuchment13, Steinberg19], and material science [Janno11, Lee25]. For instance, in medical imaging techniques such as Electrical Impedance Tomography (EIT), recovering these hidden material properties from indirect measurements enables visualization of conductivity distributions inside the human body, thereby enhancing diagnostic capabilities [Bar21, Cheney99, Patterson05]. Similarly, in geophysical contexts, inferring subsurface properties such as permeability from surface measurements is critical for resource exploration and environmental monitoring [Ewing91, Carrera88].
In recent years, machine learning techniques have been incorporated into the study of inverse problems, offering new perspectives beyond traditional approaches. Pioneering works have demonstrated that deep neural networks can solve PDE-related inverse problems directly, and improve reconstruction compared to classical methods [Jin17, Adler18]. These developments have been systematically reviewed and categorized in comprehensive surveys on deep learning for inverse problems [Arridge19, Ongie20]. Building on these advances, the neural operator framework was recently introduced to approximate solution operators between function spaces [Kovachki23], enabling the efficient resolution of parameterized PDE inverse problems. In parallel, the Finite Element Methods (FEM) has long played a central role in inverse problems, particularly in elasticity imaging. Early studies used the FEM to reconstruct material parameters from displacement data, and adjoint-based formulations later improved efficiency for biomedical applications [Maniatty89, Oberai03, Doyley12]. FEM has also been applied to other inverse settings such as optical tomography [Arridge09], and recent work has extended its use to nonlinear constitutive models and hybrid approaches that integrate FEM with neural networks for efficient PDE inversion [Xu25, Meethal23].
In this paper, we focus on the inverse problem for elliptic PDEs with variable, potentially discontinuous coefficients. The main novelty of our approach is the utilization of the Dirichlet-to-Neumann (DtN) map—a mathematical construct that encapsulates the relationship between prescribed boundary inputs and their resulting flux responses [Feldman19]. This mapping framework is especially advantageous because it allows for the reconstruction of the interior characteristics of a domain using solely boundary measurements. In scenarios such as EIT or subsurface exploration, where sensors are predominantly deployed on the boundary, this capability eliminates the need for invasive internal data collection while still providing detailed insight into the material properties within the domain. See Figure 1 for the illustration.
@@
The proposed computational framework further distinguishes itself by integrating a self-supervised machine learning algorithm [Subramanian23] with FEM used in the inner loop for solving the forward problem. This combination not only achieves high accuracy in modeling the physical processes but also enables iterative refinement of the coefficient reconstructions. The carefully designed loss functions, tailored to handle discontinuities effectively, enhance the overall fidelity of the inverse solution, making our approach robust and versatile. Ultimately, this framework offers a precise and reliable estimation of critical parameters, paving the way for significant advancements in applications ranging from non-invasive medical diagnostics to environmental and subsurface monitoring.
In summary, our novelties include
  1. 1.
    Reconstruction of an unknown coefficient using only boundary measurements, without relying on interior data.
  2. 2.
    The framework demonstrates good performance even with partial boundary observations, reflecting realistic sensing scenarios.
  3. 3.
    It can reconstruct discontinuous coefficients by incorporating carefully designed loss functions.
  4. 4.
    A self-supervised framework is employed that learns from Cauchy data pairs, without requiring labeled interior coefficients.
  5. 5.
    The proposed method integrates the Finite Element Method (FEM) into the inner loop with adjoint-based gradients, utilizing FEM as a physics-grounded solver to ensure accuracy and consistency with the underlying PDE while maintaining computational efficiency in the inversion process.
The rest of the paper is organized as follows. First, the governing system and the DtN mapping are introduced in Section 2. The overall computational framework, including the loss functions, is discussed in Section 3. Finally, in Section 4, several numerical experiments to illustrate the capabilities and the performance of our algorithm are presented.
@@ 2 Governing System
In this paper, we consider the following elliptic PDE
@@
where $\Omega$ is a bounded domain in $\mathbb{R}^{2}$, $p:\overline{\Omega}\rightarrow\mathbb{R}$ is the solution function, $f$ is the source function, $g\in H^{1/2}(\partial\Omega)$ is the given Dirichlet boundary condition, and $k$ is a strictly positive coefficient defined in $\overline{\Omega}$. Here, ${\mathbf{x}}=(x,y)$ denotes the spatial coordinate in $\mathbb{R}^{2}$.
We consider an inverse problem associated with (1), where the goal is to reconstruct the coefficient $k$ from the boundary measurements. To achieve this, we employ deep learning techniques. A straightforward approach would be to use supervised learning, which requires access to exact coefficient values of $k({\mathbf{x}})$ for all ${\mathbf{x}}\in\overline{\Omega}$. However, in most cases, obtaining complete knowledge of $k$ in $\overline{\Omega}$ is infeasible. Thus, we propose a self-supervised learning framework that does not rely on exact coefficient values of $k$ but instead leverages available data such as Dirichlet and Neumann boundary conditions to infer the coefficient.
@@ 2.1 Dirichlet-to-Neumann Map
In this subsection, we introduce the DtN map, which is central to the reconstruction of $k$. This map is defined as follows:
@@
The reconstruction of the coefficient $k$ using the DtN map has been extensively studied in the context of EIT and the Calderón problem [Feldman19]. In particular, the case when $f=0$ has been comprehensively surveyed, including the uniqueness of $k$ under partially observed boundary data [Uhlmann14].
Here, we briefly recapitulate the theorems that guarantee the existence and the uniqueness of $k$ with given conditions. Let $\Gamma\subset\partial\Omega$ be a non-empty open subset of the boundary. Then the following theorem proves the uniqueness of $k$ in the case where $f=0$.
Theorem A (Corollary 1.19 in [Uhlmann14])
@@
$\Lambda_{k_{1}}(g)=\Lambda_{k_{2}}(g)\text{ on }\Gamma\text{ for all }g\in H^{1/2}(\partial\Omega),\,\operatorname{supp}g\subset\Gamma.$
@@
Next, in the case where $f\neq 0$, we note that the uniqueness of $k$ in the entire domain $\Omega$ is not guaranteed. However, its uniqueness on the boundary $\partial\Omega$ has been established in [Isakov06] as the following theorem:
Theorem B (Theorem 5.1.1 in [Isakov06])
@@
 $\text{ If }\Lambda_{k_{1}}=\Lambda_{k_{2}},\text{ then }\partial^{\alpha}k_{1}=\partial^{\alpha}k_{2}\text{ on }\partial\Omega\cap V\text{ when }|\alpha|\leq m.$
If $k_{1},k_{2}\in C(\Omega)$ are two coefficients that are analytic in $\Omega$ and $\Lambda_{k_{1}}=\Lambda_{k_{2}}$, then $k_{1}=k_{2}$ on $\partial\Omega$ and by Strong Maximum Principle [evans10, Chapter 2], we have $k_{1}=k_{2}$ in $\Omega$.
Let $\omega$ be either $\Gamma$ or $\partial\Omega$. We refer to the case $\omega=\Gamma$ as the partially observed setting, and $\omega=\partial\Omega$ as the fully observed setting. Instead of directly using the DtN map, which is usually difficult to obtain in real-world scenarios, we utilize Cauchy data—pairs of Dirichlet and Neumann boundary measurements.
$\left\{(g,h_{g}):g\in H^{1/2}(\partial\Omega),\,\operatorname{supp}g\subset\omega,h_{g}=k\partial_{\nu}p_{g}|_{\omega}\right\}.$
These pairs serve as practical substitutes for the DtN map in our setup. In practice, it is not feasible to train on all functions $g\in H^{1/2}(\partial\Omega)$ with supp $g\subset\omega$. We consider the following problem:
Problem. Given a sufficiently large finite collection of Cauchy data:
$\left\{(g_{i},h_{g_{i}})_{i=1,\dots,N}:g_{i}\in H^{1/2}(\partial\Omega),\,\operatorname{supp}g_{i}\subset\omega,h_{g_{i}}=k\partial_{\nu}p_{g_{i}}|_{\omega}\right\},$
@@
Remark 1
The above theorems establish the theoretical uniqueness of the coefficient $k$ under certain conditions on the DtN map. Based on these results, our work reconstructs $k$ from finitely many Cauchy data, as utilizing the full DtN map is not feasible in practice. Moreover, our framework handles both the fully observed setting ($\omega=\partial\Omega$) and the partially observed setting ($\omega=\Gamma$).
Rather than directly using the values of $k$, we adopt a self-supervised learning framework where the model learns to reconstruct $k$ from Cauchy data alone [Hwang23, Hwang24]. This approach is particularly useful in practical setups where $k$ is not directly measurable.
@@ 3 Computational Algorithm
In this section, we present the numerical approach for solving Problem. First, in Subsection 3.1, we provide an overview of the proposed framework. Then, we introduce the model for reconstructing the coefficient $k$ in Subsection 3.2. Next, in Subsection 3.3, we briefly describe how to solve the forward problem using the FEM. Finally, in Subsection 3.4, we define the loss function used to train the model and describe the method for calculating its gradient to update the model parameters.
@@ 3.1 Overall Framework
Input $\bullet$ Domain points  $\overline{\Omega}$ $\bullet$ Neumann data $h_{g}=k\partial_{\nu}p_{g}|_{\omega}$ Neural Network $\mathcal{N}$ $\bullet$ Multilayer Perceptron (MLP) : Computation results: Data flow Loss Function $\mathcal{L}\bigl(h_{g},\tilde{h}_{g}\bigr)=\bigl\|h_{g}(\cdot)-\tilde{h}_{g}(\cdot)\bigr\|^{2}_{L^{2}(\omega)}$ Output $\bullet$ Reconstruct Coefficient $\{\tilde{k}({\mathbf{x}})\,|\,{\mathbf{x}}\!\!\in\!\!\overline{\Omega}\}$ Predicted Neumann $\tilde{h}_{g}=\tilde{k}\,\partial_{\nu}\tilde{p}_{g}|_{\omega}$ Solving for the PDE (FEniCS) Input $\bullet$ Source function $f$ $\bullet$ Dirichlet condition $g$ $\tilde{p}_{g}({\mathbf{x}})=\sum_{j=1}^{M}p_{j}\,\phi_{j}({\mathbf{x}})$ Parameter Updates Deep Learning Finite Element Method
@@
Figure 2 illustrates the process of reconstructing the coefficient $k$ using a neural network and FEM. The framework proceeds as follows: First, the domain points $\overline{\Omega}$ are the input to the neural network $\mathcal{N}$, which outputs the reconstruction $\tilde{k}$ of $k$. The neural network $\mathcal{N}$ is designed as a Multilayer Perceptron (MLP) that reconstructs $k$ from the input domain points. Next, the reconstructed coefficient $\tilde{k}$, along with the Dirichlet condition $g$ and source function $f$, is used to solve (1). The resulting predicted solution is denoted as $\tilde{p}_{g}$, which is represented as a linear combination of basis functions $\{\phi_{i}\}_{i=1}^{M}$ in the finite-dimensional subspace $V_{h}$. These basis functions and the finite-dimensional subspace are further detailed in Subsection 3.3. Using $\tilde{k}$ and $\tilde{p}_{g}$, the predicted Neumann data $\tilde{h}_{g}=\tilde{k}\partial_{\nu}\tilde{p}_{g}|_{\omega}$ is computed. This predicted Neumann data is then compared with the given Neumann data $h_{g}$ to calculate the loss function. Finally, the model parameters of the neural network $\mathcal{N}$ are updated to minimize the loss function. By iterating through this process, the neural network $\mathcal{N}$ learns to reconstruct $k$ more accurately.
@@ 3.2 Model for Reconstructing the Coefficient $k$
In this subsection, we propose a model for reconstructing the coefficient $k$. The model is based on a MLP architecture, which leverages the universal approximation theorem [Cybenko89] as its theoretical foundation. This theorem ensures that neural networks, including MLPs, can approximate any continuous function to a desired level of accuracy.
The neural network $\mathcal{N}$ takes the spatial variable $\mathbf{x}$} as input and reconstructs the coefficient $k$ as $\tilde{k}$. Specifically, the reconstruction is expressed as $\tilde{k}({\mathbf{x}})=\mathcal{N}({\mathbf{x}};\boldsymbol{\theta})$, where $\boldsymbol{\theta}$ represents the learnable parameters of the model, such as weights and biases. The network consists of three hidden layers, each containing 50 hidden neurons. To enable the model to capture complex natural signals and their derivatives, we use the sine function as the activation function in the hidden layers. This periodic nonlinearity allows the model to represent fine details and high-frequency patterns more effectively than traditional activation functions [Sitzmann20].
3.3 Finite Element Method for the Forward Problem
In this subsection, we briefly describe the forward problem for the solution $p$ of (1) by utilizing FEM. Here, we employ the linear continuous Galerkin method, and it is implemented by FEniCS [Logg12]. Given Dirichlet condition $g$ and coefficient $k$, (1) can be reformulated into a weak form by integration and then approximated in a finite-dimensional space.
In particular, let $V_{h}$ be a finite-dimensional subspace of $H^{1}_{0}(\Omega)$ (to simplify the presentation, we assume $g=0$), and assume $p_{h},v_{h}\in V_{h}$, with $p_{h}$ representing the approximate solution and $v_{h}$ being the test function used in the weak formulation. Using the basis functions $\{\phi_{i}\}_{i=1}^{M}$ of $V_{h}$, the weak form can be approximated as:
$\int_{\Omega}k\nabla p_{h}\cdot\nabla v_{h}\,{\rm d}{\mathbf{x}}=\int_{\Omega}fv_{h}\,{\rm d}{\mathbf{x}},\ \forall v_{h}\in V_{h}.$
Assuming $p_{h}=\sum_{i=1}^{M}P_{i}\phi_{i}$, where $P_{i}$ are constant coefficients, and substituting $v_{h}=\phi_{j}$, the resulting system becomes
$\sum_{i=1}^{M}P_{i}\int_{\Omega}k\nabla\phi_{i}\cdot\nabla\phi_{j}\,{\rm d}{\mathbf{x}}=\int_{\Omega}f\phi_{j}\,{\rm d}{\mathbf{x}},\ \forall j=1,\cdots,M.$